OpenStudy (anonymous):

For each line below, enter if the line is vertical, enter if the line is horizontal, or enter if the line is slanted (neither vertical nor horizontal). 7y=4 -4y=5x 5x+8y+3=0 5x+3=0 9x=4 8y+3=0

OpenStudy (anonymous):

7y=4 vertical -4y=5x slanted 5x+8y+3=0 slanted 5x+3=0 horizontal 9x=4 horizontal 8y+3=0 vertical

OpenStudy (anonymous):

i treid that and it says its wrong :(

OpenStudy (anonymous):

7y=4, then y = 4/7, so for every x, y stays the same: 4/7. So this line is horizontal. Can you do the next one?

OpenStudy (anonymous):

i got it thanks